OpenStudy (anonymous):

simplify radical square root of 600 ill give a metal ASAP

OpenStudy (goldeverything):

77.45966692

OpenStudy (anonymous):

\[\sqrt{100*6}= 10\sqrt{6} \]

OpenStudy (anonymous):

thnx luis